# Beyond Accuracy-Fairness: Stop evaluating bias mitigation methods solely on between-group metrics

### v1

@@ Abstract
Artificial Intelligence (AI) finds widespread applications across various domains, sparking concerns about fairness in its deployment. While fairness in AI remains a central concern, the prevailing discourse often emphasizes outcome-based metrics without a nuanced consideration of the differential impacts within subgroups. Bias mitigation techniques do not only affect the ranking of pairs of instances across sensitive groups, but often also significantly affect the ranking of instances within these groups. Such changes are hard to explain and raise concerns regarding the validity of the intervention. Unfortunately, these effects largely remain under the radar in the accuracy-fairness evaluation framework that is usually applied. This paper challenges the prevailing metrics for assessing bias mitigation techniques, arguing that they do not take into account the changes within-groups and that the resulting prediction labels fall short of reflecting real-world scenarios. We propose a paradigm shift: initially, we should focus on generating the most precise ranking for each subgroup. Following this, individuals should be chosen from these rankings to meet both fairness standards and practical considerations.
@@ 1 Introduction
In the rapidly evolving landscape of Artificial Intelligence (AI) and machine learning, the pursuit of fairness in algorithmic decision-making has emerged as a central concern. As the influence and scope of the decisions made by AI systems are increasing, there are growing concerns that the models making these decisions might unintentionally encode and even amplify human bias (Corbett-Davies and Goel, 2018). The term bias is used to describe the situation in which sensitive groups are substantially disadvantaged by an algorithm or model. One of the ways bias can seep into a model is when it is trained on biased data, following the famous garbage in, garbage out principle which emphasizes how flawed input data will result in flawed output (Geiger et al., 2020). Examples of bias in AI models are everywhere, with cases in almost every domain. In the context of hiring, a well-known use case is that of an automated Amazon recruitment system that had to be pulled because it was biased against female applicants (Dastin, 2022). Much earlier already, St George’s Hospital Medical School’s Commission for Racial Equality discovered that a computer program used for initial screenings of applicants “written after careful analysis of the way in which the staff were making these choices” unfairly rejected women and individuals with non-European sounding names (Johnson, 2021; Lowry and Macpherson, 1988). There is an abundance of examples akin to these ones.
@@
Many bias detection and mitigation methods exist, but most of them focus on "between-group fairness" where the primary objective is to rectify disparities in model predictions between distinct demographic groups. This endeavor is undeniably critical, as it aims to rectify long-standing inequalities. However, it is equally essential to acknowledge and scrutinize the complexities that exist "within" these groups, giving rise to a concept that is commonly referred to as "within-group fairness". Speicher et al. (2018) already note that many approaches to group fairness tackle only between-group issues, worsening within-group fairness as a consequence. Krco et al. (2023) highlight that the blind optimization of commonly used fairness metrics does not show who is impacted within each group, while Mittelstadt et al. (2023) emphasize that many of the currently used bias mitigation methods can make every group worse off. These issues illustrate why solely looking at fairness by measuring disparities between groups is not adequate.
While benchmarking studies attempt to gauge the performance of various bias mitigation methods, they often fall short, comparing what can be described as ‘apples to oranges’. This issue arises because different bias mitigation methods can significantly vary the number of positive instances, and by comparing them as-is, the strict constraints faced by practitioners are not taken into account. In almost every application of machine learning, such constraints are a reality. For example, banks in credit scoring are limited by an annual loan quota. Universities have a cap on student admissions each year. In cancer screening, healthcare authorities can only screen a certain percentage of the population due to capacity constraints. Similarly, in predicting recidivism, the number of available prison places limits the potential positive rate.
@@
The current state-of-art in benchmark studies of bias mitigation methods is to compare these methods on metrics based on the prediction label (such as accuracy or statistical parity difference) (Chen et al., 2023; Hort et al., 2022; Hufthammer et al., 2020; Reddy, 2022; Janssen and Sadowski, 2021; Menon and Williamson, 2018). However, the current metrics fall short in acknowledging the inherent differences in the situations they compare, rendering them insufficient for determining the superior performance of one method over another. More importantly, they fail to take into account the constraints in positive rate that are faced in the real world. For example, our results in Section 4 demonstrate that for the Adult Income dataset one bias mitigation method yields a positive rate of 0.5% while another results in a positive rate of 39.3%, whereas the actual positive rate of the dataset is 23.9%. If we compare the accuracy of these two mitigation methods, we are comparing two entirely distinct points of the data distribution, both of which may be inapplicable in real-world settings. Moreover, the majority of bias mitigation methods do not result in the satisfaction of the fairness metrics, necessitating additional post-hoc interventions after their application.
We argue that the focus should shift to fairness within groups, and the crucial question to be answered should be: how can we select the best individuals within each group? Machine learning models not only return prediction labels, they also return prediction scores, which creates an inherent ranking of instances. Given that the prediction labels from many mitigation methods are often impractical due to their unrealistic positive rates and non-compliance with the selected fairness metric, it is more logical to concentrate on the prediction scores produced by each mitigation method. The objective should be to have the machine learning model rank each subgroup as accurately as possible, a metric conveniently measured by the Area Under the Curve (AUC). Once the most accurate ranking is established, post-hoc fairness methods can be employed to select the highest-ranked individuals within each group, aligning with both industry constraints and fairness metrics. [^1] This framework operates under the assumption of having access to the sensitive attribute(s).
Figure 1: The ranking within each group is shifted by most preprocessing and inprocessing bias mitigation methods. The current state-of-the-art is to measure the accuracy and fairness metrics at one specific classification threshold.
When striving to establish the most optimal ranking, it is crucial to differentiate between pre- and inprocessing bias mitigation methods on the one hand, and postprocessing methods on the other. Only pre- and inprocessing methods have the capacity to alter this ranking by either fine-tuning the training data or incorporating fairness constraints during model training, as this will change the prediction scores of the machine learning model. In contrast, postprocessing methods merely flip prediction labels based on the scores of the initial machine learning model, without altering the inherent ranking of instances. Figure 1 visualizes how the inherent ranking is shifted by pre- and inprocessing mitigation methods, emphasizing the prevailing practice of assessing performance and fairness metrics at a single classification threshold. Furthermore, it underscores the tendency to overlook the impact of fairness interventions on the rankings within each subgroup.
The contributions of this paper are threefold: Firstly, we provide insights into the operational dynamics of bias mitigation methods and illustrate how some methods will significantly impact the ranking within each group, while others will not. Secondly, we criticize the current approach to comparing bias mitigation methods, as it compares situations that are significantly distinct and not applicable in real-world settings. Therefore, we propose a paradigm shift, advocating for a transition from focusing on measuring the disparity in prediction labels between-groups to establishing the best prediction scores within-groups. This leads us to our last contribution: a novel framework. This framework consists of a two-step approach: first, optimizing for the most accurate ranking within each subgroup, and then, selecting the highest-ranked individuals from each subgroup, aligning with fairness metrics and industry constraints.
@@ 2 Background
Before we go further, it is important to define some of the key terminology that is often used in the fairness literature. A sensitive attribute refers to a characteristic or feature(s) of an individual that are considered sensitive, often with respect to potential discrimination. This can include attributes such as race, gender, age, religion, sexual orientation, or any other factor that could be the basis of unfair treatment. Consequently, a protected group typically refers to the demographic group that is at risk of being unfairly treated or discriminated against based on their sensitive attribute, while the privileged group is the demographic category that is typically not subject to unfair treatment based on that sensitive attribute. In this paper, we operate under the assumption of a single binary sensitive attribute, implying the existence of a protected group and a privileged group. Fairness metrics are quantitative measures used to assess the fairness of AI models, while fairness (or bias) mitigation strategies are techniques and approaches used to modify AI models to reduce bias or discrimination.
@@
In the computer science community, a plethora of fairness metrics have been proposed (Corbett-Davies and Goel, 2018). One of the most popular approaches is the group fairness metric, which quantifies the fairness of a machine learning model across different demographic or sensitive groups, aiming to identify disparities in the outcome between these groups. One of the simplest and most commonly used definitions is demographic parity (or statistical parity), which states that the positive classification rate must be the same regardless of the value of the protected attribute. In our example of hiring, this means that a model must invite equal percentages of white and black applicants for an interview (if race is the sensitive attribute) or of male and female applicants (if gender is the sensitive attribute). Other commonly used metrics include equalized opportunity, which states that there should be an equal proportion of true positives in both groups, and equalized odds, which examines whether both the proportion of true positives and true negatives is equal across groups. Besides these, many other fairness metrics exist, and the issue is that most of them cannot be optimized at the same time (Kleinberg et al., 2016). Deciding upon a group fairness metric to optimize thus means already imposing a certain world view.
@@
A common starting point for designing a fair algorithm is simply to exclude sensitive attributes from the model, which is known as fairness through unawareness. The limitations of fairness through unawareness have been commonly addressed, with the most fundamental limitation being ‘the proxy problem’. The proxy problem states that the omission of sensitive attributes can lead the machine learning model to rely on proxy variables that indirectly encode the information contained in the sensitive attribute and hence still introduce bias into the model’s decision-making process. A classic example of the proxy problem is the use of zip codes in the United States as a proxy for racial information, as these two attributes tend to be heavily correlated. This has prompted many to argue that proxies should be excluded from the dataset as well, however, this is very difficult to operationalize (Corbett-Davies and Goel, 2018). This is because every attribute used in the machine learning model can be at least partially correlated with the sensitive attribute; and often, even strongly correlated attributes may be considered legitimate factors on which to base decisions (for example, education in the case of hiring) (Corbett-Davies and Goel, 2018).
@@
This illustrates that creating a fair machine learning model is a tedious process. In response, many bias mitigation methods that claim to improve fairness, have been introduced. We can divide most of them into 3 categories: preprocessing, inprocessing and postprocessing. The idea behind preprocessing methods is that they will change the representation of the data before the machine learning model is learned. Inprocessing methods improve fairness during the training process, while postprocessing methods attempt to modify the posteriors in a way that satisfies fairness constraints. These bias mitigation methods focus on satisfying the aforementioned group fairness metrics that measure disparities between groups (Chen et al., 2023).

### v3

# Reranking individuals: The effect of fair classification within-groups

@@ Abstract
Artificial Intelligence (AI) finds widespread application across various domains, but it sparks concerns about fairness in its deployment. The prevailing discourse in classification often emphasizes outcome-based metrics comparing sensitive subgroups without consideration of the differential impacts within subgroups. Bias mitigation techniques do not only affect the ranking of pairs of instances across sensitive groups, but often also significantly affect the ranking of instances within these groups. Such changes are hard to explain and raise concerns regarding the validity of the intervention. Unfortunately, these effects remain under the radar in the accuracy-fairness evaluation framework that is usually applied. In this paper, we illustrate the effect of several popular bias mitigation methods, and how their output often does not reflect real-world scenarios.
@@ 1 Introduction
In the rapidly evolving landscape of Artificial Intelligence (AI) and machine learning, the pursuit of fairness in algorithmic decision-making has emerged as a central concern. As the influence and scope of the decisions made by AI systems are increasing, there are growing concerns that the models making these decisions might unintentionally encode and even amplify human bias (Corbett-Davies and Goel, 2018). Algorithmic bias describes situations where sensitive groups are substantially disadvantaged by an algorithm or model. One of the ways bias can seep into a model is when it is trained on biased data, following the famous garbage in, garbage out principle which emphasizing that flawed input data results in flawed output (Geiger et al., 2020). Examples of biased AI models are everywhere, with cases in almost every domain. In the context of hiring, a well-known case is that of an automated Amazon recruitment system that had to be pulled because it was biased against female applicants Dastin (2022). Much earlier already, St George’s Hospital Medical School’s Commission for Racial Equality discovered that a computer program used for initial screenings of applicants “written after careful analysis of the way in which the staff were making these choices” unfairly rejected women and individuals with non-European sounding names Johnson (2021); Lowry and Macpherson (1988). There is an abundance of examples akin to these ones.
@@
In this paper, we focus on fair classification, which aims to ensure that algorithms make unbiased decisions across groups. Many bias detection and mitigation methods exist, but most of them focus on “between-group fairness” where the main goal is to rectify disparities in model predictions between distinct demographic groups. This is undeniably critical, as it aims to rectify long-standing inequalities. However, it is also important to consider the complexities that exist “within” these groups, giving rise to a concept that is commonly referred to as “within-group fairness”. Speicher et al. (2018) already note that many approaches to group fairness tackle only between-group issues, worsening within-group fairness as a consequence. Krco et al. (2023) highlight that the blind optimization of commonly used fairness metrics does not show who is impacted within each group, while Mittelstadt et al. (2023) emphasize that many of the currently used bias mitigation methods can make every group worse off. These issues illustrate why solely looking at fairness by measuring disparities between groups is not adequate.
While various benchmarking studies attempt to evaluate the performance of bias mitigation methods, they often fall short, comparing what can be described as ‘apples to oranges’. This issue arises because different bias mitigation methods can significantly vary the number of positive instances, and by comparing them as-is, the actual situation faced by practitioners is not taken into account. Scantamburlo et al. (2024) argue that the ultimate decision of an automated system is informed by the prediction model, but in nearly all cases is also influenced by additional parameters such as quota or business rules. They make a distinction between the prediction model and the decision-making system, and discuss how the field of fair machine learning tends to blur the boundaries between the two concepts (Scantamburlo et al., 2024). Kwegyir-Aggrey et al. (2023) confirm that when deploying a classifier in the real world, practitioners typically need to tinker with the threshold to make sure the model predictions meet their domain-specific needs.
@@
This is why directly comparing these methods on prediction labels as-is is not a good idea, yet this approach remains common in current benchmarking studies (Chen et al., 2023; Hort et al., 2022; Hufthammer et al., 2020; Reddy, 2022; Janssen and Sadowski, 2021; Menon and Williamson, 2018). The existing metrics fall short in acknowledging the inherent differences in the situations they compare, rendering them insufficient for determining the superior performance of one method over another. More importantly, they fail to take into account the constraints in positive decision rate that are faced in the real world. For example, our results demonstrate that for the Adult Income dataset one bias mitigation method yields a positive decision rate of 0.5% while another results in a positive decision rate of 39.3%, whereas the actual positive rate of the dataset is 23.9%. This illustrates that the process of enforcing fairness can significantly alter the overall selection rate of the model. If we would compare the accuracy of these two mitigation methods, we are actually examining two entirely distinct points of the ROC curve, both of which may be inapplicable in real-world settings. Moreover, the majority of bias mitigation methods do not result in the satisfaction of the fairness metrics, necessitating additional post-hoc interventions after their application. To address these issues, one should separate the function of the prediction model from the decision-making context (Scantamburlo et al., 2024). This is necessary to distinguish between performance changes that are solely due to the model selecting more or less individuals, and those caused by the model selecting ‘worse’ individuals.
Following this, we explore the distinct impact that each bias mitigation method has. These methods attempt to make the model output more fair by imposing constraints at the group level, but they also have a significant effect within each group. While the primary goal is to increase the selection of individuals from the protected group and reduce selections from privileged groups, these methods can also change which specific individuals are selected within each group.  [^1] In some cases, due to changes in the global selection rate, fewer individuals from a protected group may be selected while still achieving greater overall fairness—an outcome that is far from ideal. This seemingly arbitrary reshuffling within each group, as a side effect of fair classification, is currently not studied. In this process, certain individuals (from both the protected and the privileged group) who were initially assigned a positive prediction label will now be labelled negatively, while some group members that initially received a negative label will be switched to a positive one. This reranking process within each group is not per definition negative -as we will discuss- but it deserves more attention. Machine learning models not only return prediction labels, they also return prediction scores, which creates an inherent ranking of instances. We will compare the rankings generated by each bias mitigation method to analyze the effect within subgroups, independent of changes in the positive decision rate.
In summary, even though the primary purpose of using bias mitigation methods is typically to boost selection form protected group, these methods often yield unintended consequences. In this paper, we discuss two of these consequences:
  • The global selection rate is often modified by the bias mitigation methods, which makes the resulting prediction labels not comparable and not applicable in real-world settings. We demonstrate this in Section 5.2, and argue that no distinction is made between the prediction model and the decision-making context.
  • The bias mitigation methods will not only change who is selected between the groups, but also who is selected within each group. We illustrate for several methods and datasets how some methods will shift the ranking within each group in Section 5.1. We discuss the theory behind this in Section 3.
@@ 2 Background
Before we go further, it is important to define some of the key terminology that is often used in the fairness literature. A sensitive attribute refers to a characteristic or feature of an individual that is considered sensitive, often with respect to potential discrimination. This can include attributes such as race, gender, age, religion, sexual orientation, or any other factor that could be the basis of unfair treatment. Consequently, a protected group typically refers to the demographic group that is at risk of being unfairly treated or discriminated against based on their sensitive attribute, while the privileged group is the demographic category that is typically not subject to unfair treatment based on that sensitive attribute. In this paper, we operate under the assumption of a single binary sensitive attribute, implying the existence of a protected group and a privileged group. Fairness metrics are quantitative measures used to assess the fairness of AI models, while fairness (or bias) mitigation strategies are techniques used to either learn an AI model that is fair by design or modify AI models to reduce bias.
@@
In the computer science community, a plethora of fairness metrics have been proposed (Corbett-Davies and Goel, 2018). One of the most popular approaches is the group fairness metric, which quantifies the fairness of a machine learning model across different demographic or sensitive groups, aiming to identify disparities in the outcome between these groups. One of the simplest and most commonly used definitions in this category is demographic parity (or statistical parity), which states that the positive decision rate must be the same regardless of the value of the protected attribute. In our example of hiring, this means that a model must invite equal percentages of white and black applicants for an interview (if race is the sensitive attribute) or of male and female applicants (if gender is the sensitive attribute). Other commonly used metrics include equalized opportunity, which states that there should be an equal proportion of true positives in both groups, and equalized odds, which examines whether both the proportion of true positives and true negatives is equal across groups. Besides these, many other fairness metrics exist, and the issue is that most of them cannot be optimized at the same time Kleinberg et al. (2016). Deciding upon a group fairness metric to optimize thus means already imposing a certain world view. Another popular approach to assess fairness in machine learning models is individual fairness, which demands that similar individuals receive similar outcomes in a decision-making process, regardless of their group membership (Dwork et al., 2012; Binns, 2020). Dwork et al. (2012) argue that instead of focusing on a group, we tend to care more about the individual. This notion is related to the contributions of this paper, as we will also argue that satisfying group fairness metrics is not necessarily fair from the viewpoint of the individual.
@@
A common starting point for designing a fair algorithm is simply to exclude sensitive attributes from the model. However, the limitations of this approach have been commonly addressed, with the most fundamental limitation being the proxy problem (Prince and Schwarcz, 2019). The proxy problem states that the omission of sensitive attributes can lead the machine learning model to rely on proxy variables that indirectly encode the information contained in the sensitive attribute and hence still introduce bias into the model’s decision-making process. A classic example of the proxy problem is the use of zip codes in the United States as a proxy for racial information, as these two attributes tend to be heavily correlated. This has prompted many to argue that proxies should be excluded from the dataset as well, however, this is very difficult to operationalize (Corbett-Davies and Goel, 2018). This is because every attribute used in the machine learning model can be at least partially correlated with the sensitive attribute; and often, even strongly correlated attributes may be considered legitimate factors on which to base decisions (for example, education in the case of hiring) (Corbett-Davies and Goel, 2018).
@@
This illustrates that creating a fair machine learning model is a tedious process. In response, many bias mitigation methods that claim to improve fairness, have been introduced. We can divide most of them into three categories: preprocessing, inprocessing and postprocessing. Each category targets a different stage of the machine learning pipeline to ensure fairness. The idea behind preprocessing methods is that they will change the representation of the data before the machine learning model is learned, and as such neutralize any prejudiced information that could affect the model’s decision. Inprocessing methods improve fairness during the training process, by incorporating fairness constraints in the learning algorithm and striving for a balance between accuracy and fairness. Postprocessing methods intervene after the model has made its predictions, by adjusting the outcomes to satisfy fairness criteria. We will discuss the used bias mitigation methods in more detail in the Materials and Methods section. These bias mitigation methods focus on satisfying the aforementioned group fairness metrics that measure disparities between groups Chen et al. (2023).
Another research area that we should consider is the area of fair ranking. Yang and Stoyanovich (2017) measure whether a ranking is fair by comparing the distribution of protected and non-protected candidates on different prefixes of the list, while Zehlike et al. (2017) illustrates how group fairness metrics can be satisfied for different prefixes of the ranking. Yang et al. (2019) demonstrate how adding diversity constraints to ranking algorithms can reduce in-group fairness, a concept related to our measure of within-group fairness. However, it is crucial to highlight the differences with this study, as we study fair classification, which focuses on equal outcomes between groups, and which does not entail that the ranking distribution should be fair.  [^2] This can be fixed by postprocessing methods, for example by using different thresholds for each group, for the required positive decision rate. Fair ranking will be more strict, as it requires that the fairness metrics are satisfied for different prefixes of the ranked list. We will show that several methods that are used for fair classification have as side effect that they change the rankings within each group, which in turn has influence on the final prediction labels of each individual. This seemingly arbitrary ‘reshuffling’ as a consequence of fair classification is currently not studied and deserves more attention.
3 Theory
This section offers a theoretical foundation for analyzing how bias mitigation methods impact group-specific rankings. Such a framework highlights the importance of auditing fairness algorithms and explains why measuring differences in rankings provides meaningful insights into a model’s behavior. The key takeaway from this section is that bias mitigation methods have two potential outcomes: they either address within-group bias or must align theoretically with Threshold Optimization to be optimal. A more comprehensive explanation of the theory underlying this section is available in Favier and Calders (2024).
Consider a space of individuals from whom we have collected attributes $X=(X_{1},\dots,X_{n})$, as well as an assumed-binary sensitive attribute $A$, and a binary label $Y\in\{0,1\}$.
Every individual has an unbiased and fair probability of receiving the positive label $p(Y=1|X=x,A=a)$. At the same time, we have access to a biased model, trained on biased data, that provides a score $S(x,a)\in[0,1]$, which can be used as a biased proxy for $p(Y=1|X=x,A=a)$.
More specifically, $S(x,a)$ is assumed to be the best fairness-agnostic model achievable, capable of providing the most accurate score based on biased data. In this sense, $S(x,a)$ accurately reflects the unfair probability distribution captured in the collected data. In other words, we will assume that $S(x,a)$ has no epistemic uncertainty, and any deviation of $S(x,a)$ from the fair $p(Y=1|X=x,A=a)$ is due to bias in the data. Given a distribution on $X\times A$, both $p(Y=1|X=x,A=a)$ and $S(x,a)$ can be viewed as random variables on $[0,1]$, which we assume to be continuous.
Example 3.1 (part 1).
[…]
$p(Y=\text{\leavevmode\ltxml@oqmark@open\textquotedblleft\penalty 10000\hskip-0.0002pt\hskip 0.0002pthigh\textquotedblright\ltxml@oqmark@close{}}|X,A=\male)=p(Y=\text{\leavevmode\ltxml@oqmark@open\textquotedblleft\penalty 10000\hskip-0.0002pt\hskip 0.0002pthigh\textquotedblright\ltxml@oqmark@close{}}|X,A={\female})$
Notice that, since $X$ and $A$ are independent, fair probabilities satisfy demographic parity. However, if we were to train a model on biased data, we would have that
$S(X,A={\female})=90\%\cdot S(X,A={\male})$
This occurs because the model has learned biased probabilities, and even with calibration on the data, it will remain biased. Furthermore, the scores no longer adhere to demographic parity.
Definition 3.2.
A decision is a function $\hat{Y}\colon U\subseteq X\times A\to{0,1}$ that we can evaluate based on how well it performs on each label. In particular, for all $y\in\{0,1\}$, we define the following values:
$P_{f}(\hat{Y}=y|Y=y)$ $:=\dfrac{\int_{U}|1-y-\hat{Y}(x,a)|\cdot|1-y-p(Y=1|X=x|A=a)|\text{d}(x,a)}{\int_{U}|1-y-p(Y=1|X=x,A=a)|\text{d}(x,a)}$
$P_{u}(\hat{Y}=y|Y=y)$ $:=\dfrac{\int_{U}|1-y-\hat{Y}(x,a)|\cdot|1-y-S(x,a)|\text{d}(x,a)}{\int_{U}|1-y-S(x,a)|\text{d}(x,a)}$
where $P_{f}(\hat{Y}=y|Y=y)$ is the true positive/negative rate according to fair probabilities, and $P_{u}(\hat{Y}=y|Y=y)$ is the true positive/negative rate according to the unfair score, depending on the considered label $y$.
This allows us to define the following:
Definition 3.3 (Pareto Order).
Given two decisions $\hat{Y},\hat{Y}^{\prime}$, we will say that $\hat{Y}$ is Pareto non-inferior to $\hat{Y}^{\prime}$ according to fair probabilities, and write $\hat{Y}\geq_{f}\hat{Y}^{\prime}$ if and only if
$P_{f}(\hat{Y}=1|Y=1)\geq P_{f}(\hat{Y}^{\prime}=1|Y=1)\text{ and }P_{f}(\hat{Y}=0|Y=0)\geq P_{f}(\hat{Y}^{\prime}=0|Y=0)$
Similarly, we will say that $\hat{Y}$ is Pareto non-inferior to $\hat{Y}^{\prime}$ according to unfair probabilities, and write $\hat{Y}\geq_{u}\hat{Y}^{\prime}$ if and only if
$P_{u}(\hat{Y}=1|Y=1)\geq P_{u}(\hat{Y}^{\prime}=1|Y=1)\text{ and }P_{u}(\hat{Y}=0|Y=0)\geq P_{u}(\hat{Y}^{\prime}=0|Y=0)$
It is possible to prove that Pareto maximal decisions according to fair probabilities are all and only those of the following form almost everywhere:
$\hat{Y}(x,a)=\begin{cases}1&\text{if }p(Y=1|X=x,A=a)>\tau\\ 0&\text{otherwise}\end{cases}$
for some $\tau\in[0,1]$. Similarly, for Pareto maximal decisions according to unfair probabilities, the same property holds with respect to $S(x,a)$.
Example 3.4 (part 2).
After training a model for the previous example, we obtain the following scores $S(x,a)$:
$S(x,a)=\begin{cases}x/50&\text{if }a={\male}\\ 90\%x/50&\text{if }a={\female}\end{cases}$
So the following decision function is used to make predictions:
$\hat{Y}(x,a)=\begin{cases}1&\text{if }S(x,a)>0.5\\ 0&\text{otherwise}\end{cases}$
According to unfair probabilities, this decision is Pareto maximal. However, it is not maximal according to the fair probabilities, since according to fair probabilities the decision is equivalent to
$\hat{Y}(x,a)=\begin{cases}1&\text{if }a=\male\text{ and }p(Y=1|X=x,A=a)>0.5\\ 1&\text{if }a=\female\text{ and }p(Y=1|X=x,A=a)>0.55\overline{5}\\ 0&\text{otherwise}\end{cases}$
which depends specifically on the sensitive attribute.
Under this framework, we can prove that the absence of within-group bias is equivalent to claiming that Pareto maximal decisions on fair probabilities correspond to decision that are still Pareto maximal according to unfair probabilities but only on each sensitive group. In other words, either we believe that within group bias exists or we already know the optimal approach to make fair decisions.
More formally, we formalize the assumption that every fair Pareto maximal decision corresponds to unfair Pareto maximal ones for each sensitive group as follows
Definition 3.5 (Affirmative Action Assumption).
Let $\hat{Y}\colon X\times A\to\{0,1\}$ be a maximal decision according to fair probabilities. For every $a\in A$, there exists a maximal decision according to unfair probabilities $\hat{Y}_{a}\colon X\times\{a\}\to\{0,1\}$ such that $\hat{Y}=\bigcup_{a\in A}\hat{Y}_{a}$
Example 3.6 (part 3).
In our running example we have that the Affirmative Action Assumption does hold, since the following holds:
$\hat{Y}(x,a)=\begin{cases}1&\text{if }p(Y=1|X=x,A=a)>\tau\\ 0&\text{otherwise}\end{cases}\iff\hat{Y}(x,a)=\begin{cases}1&\text{if }a=\male\wedge S(x,a)>\tau\\ 1&\text{if }a=\female\wedge S(x,a)>9\tau/10\\ 0&\text{otherwise}\end{cases}$
which means that each fair optimal decision is equivalent to a decision that utilizes two different thresholds, one for each sensitive group.
The following theorem asserts that if we have a Pareto maximal decision according to fair probabilities, then, in absence of within group bias, this decision can be decomposed into two maximal decisions - one for each sensitive group.
Theorem 3.7.
The Affirmative Action assumption holds if and only if for all $a\in A$ it holds that for almost all $(x,a),(x^{\prime},a)\in X\times\{a\}$
$p(Y=1|X=x,A=a)\leq p(Y=1|X=x^{\prime},A=a)\Rightarrow S(x,a)\leq S(x^{\prime},a)$
Proof.
The Proof of this theorem can be found in Section A. ∎
This Theorem implies that reranking within each group only makes sense when the Affirmative Action assumption does not hold, and there is some form of within-group bias. When this is the case, it makes sense that we need to make changes within each group as well. However, when the Affirmative Action assumption holds and there is no within-group bias, there is no practical reason to use any fairness intervention other than Threshold Optimization, as research has shown that for fairness measures such as Demographic Parity and Equal Opportunity, Threshold Optimization can always find an optimal solution. We will see in Section 5.1 that many of the used bias mitigation methods do lead to reranking of individuals within each group, while they do not explicitly assume any within-group bias.